can you please derive the equation for drift velocity

Grade:10

1 Answers

Umakant biswal
7 years ago
@ padma 
let the electron have  charge e , mass m , and kept in electric field inside a conducter . 
THE FORCE experienced by the electron , F = eE 
acceleration due to force , = eE/m 
now displacement s due to acceleration s = ½ at square 
average velocity v(d) 
drift velocity = total displacement / time taken = s/t 
so, by deriving 
V(d) = eE/ 2m 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..